OpenStudy (anonymous):

Solve: 49x^2 – 28x = -4

OpenStudy (anonymous):

x=2/7

OpenStudy (anonymous):

x^2 - 4/7 x + 4/49 = 0 ( x - 2/7 )( x - 2/7 ) = 0 => x = 2/7